OpenStudy (anonymous):

cot^2 135°- sin 30° + 4 tan 45° (1)^2 - 1/2 + 4(1)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The only issue I have is cot(135) is NOT equal to +1 it's actually equal to -1

jimthompson5910 (jim_thompson5910):

but it won't matter because (-1)^2 = (1)^2 = 1

jimthompson5910 (jim_thompson5910):

it's still a good idea to evaluate each trig expression correctly

OpenStudy (anonymous):

Ok Ill ask my teacher about cot 135 because she gave me the notes.. but still how does everything equal to 1?

jimthompson5910 (jim_thompson5910):

cot(135) = 1/tan(135) cot(135) = 1/(-1) cot(135) = -1

jimthompson5910 (jim_thompson5910):

evaluate (-1)^2 - 1/2 + 4(1) to get ???

OpenStudy (anonymous):

it gives me 4 1/2

jimthompson5910 (jim_thompson5910):

I'm getting 4.5 which is equivalent to \(\Large 4 \frac{1}{2}\)

jimthompson5910 (jim_thompson5910):

so you are correct

OpenStudy (anonymous):

Ok thanks ;) @jim_thompson5910